You are driving at 53 mi/h in a 30-mi/h zone. Spotting a police officer ahead, you brake at a constant rate to 27 mi/h, traveling 294 ft in so doing. What time interval do you need to reach your final speed? Use one decimal place in your answer.
Answer #1

1 mi/h = 0.44704 m/s

53 mi/h = 53*0.44704 = 23.7 m/s

27 mi/h = 27*0.44704 = 12.07 m/s

294 ft = 294*0.3048 = 89.6 m

using the equation we have

v^2 - u^2 = 2aS

12.07^2 - 23.7^2 = 2*a*89.6

a = -2.32 m/s2

v = u + at

12.07 = 23.7 - 2.32*t

t = 5.0 s